OpenStudy (anonymous):

Find the inverse of the function. f(x) = 4x^3 - 7

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

The basic idea is to replace f(x) with y, swap x and y, then solve for y to get the inverse. Similar Example (just with different numbers) \[\Large f(x) = 19x^3 + 2\] \[\Large y = 19x^3 + 2\] \[\Large x = 19y^3 + 2\] \[\Large x-2 = 19y^3\] \[\Large 19y^3 = x-2\] \[\Large y^3 = \frac{x-2}{19}\] \[\Large y = \sqrt[3]{\frac{x-2}{19}}\] \[\Large f^{-1}(x) = \sqrt[3]{\frac{x-2}{19}}\] --------------------------------------------------------------- So the inverse of \(\Large f(x) = 19x^3 + 2\) is \(\Large f^{-1}(x) = \sqrt[3]{\frac{x-2}{19}}\)

OpenStudy (anonymous):

so it would be \[\sqrt[3]{\frac{ x+7 }{ 4 }}\]

jimthompson5910 (jim_thompson5910):

correct